Hi, I need help solving this 7th grade math problem:
Write each sequence of numbers in exponent form;
8,24,72,216
Thanks!!

Answer by Edwin McCravy(20056)


I'm not sure what you are looking for.  Maybe it's this:

8 = 4×2 = 2×2×2 = 23

24 = 4×6 = 2×2×2×3 = 23×31

72 = 8×9 = 4×2×3×3 = 2×2×2×3×3 = 23×32

216 = 2×108 = 2×2×54 = 2×2×2×27 = 2×2×2×3×9 = 2×2×2×3×3×3 = 23×33

If not then maybe it's this:

8, 24, 72, 216

since we multiply by 3 each time to get the next term of the sequence:

an = 8×3n-1

If neither of these is what you want, then post again, copying every
word in the instructions so we will know what you are looking for.

Edwin